OpenStudy (anonymous):

simplify using absolute value 5^√-1/32

OpenStudy (saifoo.khan):

\[-\frac12\]

OpenStudy (anonymous):

ok so how did you get this? i would like to know how to do it not just the answer. sorry I like to learn not just find out the answer.

OpenStudy (anonymous):

saifoo! long time

OpenStudy (anonymous):

just talking about you http://www.youtube.com/watch?v=i8npC1TAIao

OpenStudy (saifoo.khan):

\[\Large \sqrt[5]{-32} \to -2\] Or in other words, \[-\frac12 \times-\frac12 \times-\frac12 \times-\frac12 \times-\frac12 ==-\frac{1}{32}\]

OpenStudy (saifoo.khan):

Lol @ sat.

OpenStudy (anonymous):

\[\left(\frac{1}{2}\right)^5=-\frac{1}{32}\]

OpenStudy (saifoo.khan):

Where's the negative sign? :O

OpenStudy (saifoo.khan):

\[\left(-\frac{1}{2}\right)^5=-\frac{1}{32}\]

OpenStudy (saifoo.khan):

\[\left(\frac{1}{2}\right)^5 \ne -\frac{1}{32}\]

OpenStudy (anonymous):

i got lost listening to the sagoo remix, of kinna sohna tenu featuring c jam blues in the middle

OpenStudy (saifoo.khan):

Haha. I figured. :D

OpenStudy (anonymous):

ty both so much I get it now

OpenStudy (saifoo.khan):

welcome! ^_^